Please assist with this equation.
Solve using the substitution method. Show your work. If the system has no solution or an infinite number of solutions, state this.
8x – 5y = 3
7x + y = 51

Answer by Greenfinch(383)

y = 51 - 7x
Therefore first equation becomes 8x - 5 (51 - 7x) = 3
So 8x - 255 + 35x = 3
or 43x = 258
x = 258/43
x = 6
Second equation becomes 7x6 + y = 51
42 + y = 51
y = 51 - 42 = 9
Check in equation 1 8x6 - 5x9 = 48 - 45 = 3